OpenStudy (anonymous):

What are the polar coordinates of (0,-sqrt10)

OpenStudy (anonymous):

As the x coordinate is 0, we see it lies on the (negative) y axis, and so theta is 3pi/2 the radius is the square root of the sum of the squares of x and y, so \[r = \sqrt{0^2 + (-\sqrt{10})^2} = \sqrt{10}\]

OpenStudy (anonymous):

The x coordinate was sqrt10 but it doesn't accept theta as 3pi/2

OpenStudy (anonymous):

does it have to be in the interval (-pi, pi) ? Then subtract 2pi from it: -pi/2 is the same angle.